# A Mixed-Integer Conic Program for the Multi-Agent Moving-Target Traveling Salesman Problem

### v1

@@ Abstract
The Moving-Target Traveling Salesman Problem (MT-TSP) aims to find a shortest path for an agent that starts at a stationary depot, visits a set of moving targets exactly once, each within one of their respective time windows, and then returns to the depot. In this paper, we introduce a new Mixed-Integer Conic Program (MICP) formulation that finds the optimum for the Multi-Agent Moving-Target Traveling Salesman Problem (MA-MT-TSP), a generalization of the MT-TSP involving multiple agents. We obtain our formulation by first restating the current state-of-the-art MICP formulation for MA-MT-TSP as a Mixed-Integer Nonlinear Nonconvex Program, and then reformulating it as a new MICP. We present computational results to demonstrate the performance of our approach. The results show that our formulation significantly outperforms the state-of-the-art, with up to a two-order-of-magnitude reduction in runtime, and up to over 90 % tighter optimality gap.
@@ I Introduction
The objective of this paper is to find exact solutions for the Multi-Agent Moving-Target Traveling Salesman Problem (MA-MT-TSP), a generalization of the MT-TSP with multiple agents. Currently, the only exact solver for this problem is the MICP introduced in [27]. Hence, we take this as our baseline. Here, each target is assumed to move along its own line, and have one associated time window. However, in this paper, we consider a less restricted case where each target traverse a path made of piecewise-linear segments, and have more than one associated time window (Fig. 1). Therefore, we first extend the baseline MICP to accommodate these changes. Then, we pose this extended formulation as a mixed-integer nonlinear nonconvex program, which we then finally reformulate as a new MICP, similar to the works in [23], and [18].
@@ III Baseline MICP for MA-MT-TSP
Although this MICP describes the MA-MT-TSP well, It is challenging to solve in practice. This is possibly due to the big-$M$ constraints (7) and (9), which may lead to poor convex relaxations, as well as numerical instabilities, as noted by the authors in [27]. In the next section, we present an alternative MICP based on the ideas in [23] and [18], that finds the optimum for the MA-MT-TSP, much faster than the baseline. Prior to presenting this formulation, we restate the baseline as a mixed-integer nonlinear nonconvex program. This will aid us in proving that an optimal solution for our new MICP indeed provides an optimal solution for the MA-MT-TSP.
@@ IV A New MICP for the MA-MT-TSP
IV-A Mixed-Integer Nonconvex Program for MA-MT-TSP
This section details how the baseline MICP for the MA-MT-TSP can be restated as a mixed-integer nonlinear nonconvex program. We refer to this program as the nonconvex formulation for simplicity. First, we define the decision variables. For each node $i\in V$ and agent $k\in V_{agt}$, we reuse the variable $t_{i,k}$ from the baseline. In addition, we also define $p_{i,k}$ as a new auxiliary variable. For each edge $e=(i,j)\in E$, we reuse the variables $y_{e,k}$, $l_{e,k}$, and $l_{e,k}^{xy}$ from the baseline. The $l_{e,k}$ variable is defined with the help of new real auxiliary variables $z_{e,k,p}$ and $z_{e,k,t}$ representing the products $y_{e,k}p_{i,k}$ and $y_{e,k}t_{i,k}$ respectively, and variables $z^{\prime}_{e,k,p}$ and $z^{\prime}_{e,k,t}$ representing $y_{e,k}p_{j,k}$ and $y_{e,k}t_{j,k}$ respectively. Observe how in the baseline, $l_{e,k}$ was defined with the help of $\overline{l}_{e,k}$ and big-$M$ constraints instead. We now present the nonconvex formulation below.
@@ IV-A Mixed-Integer Nonconvex Program for MA-MT-TSP
$\begin{split}&z_{e,k,p}=y_{e,k}p_{i,k},\;z_{e,k,t}=y_{e,k}t_{i,k},\;z^{\prime}_{e,k,p}=y_{e,k}p_{j,k},\\ &z^{\prime}_{e,k,t}=y_{e,k}t_{j,k},\;\;\forall\;e=(i,j)\in E,\;\forall\;k\in V_{agt}.\end{split}$ (16)
@@ IV-B A New MICP for MA-MT-TSP
From an optimal solution to the MICP, optimal agent tours for the nonconvex formulation can be obtained by considering the subset of edges $E_{tour}:=\{e\in E:y_{e}=1\}$ and finding all the $1\leq\alpha\leq m$ vertex-disjoint paths starting at $s$ and ending at $s^{\prime}$ within $E_{tour}$. This can be done using a graph search algorithm such as Depth-First Search (DFS). Since all the agents are homogeneous and share the same depot, any of the paths can be assigned to any agent. Let $E_{tour}^{k}\subseteq E_{tour}$ denote the path assigned to an agent $k\in V_{agt}$. Note that $E_{tour}=\bigcup_{k\in V_{agt}}E_{tour}^{k}$. Also note that it is possible $E_{tour}^{k}=\emptyset$ for some $k$ in the case where some of the agents are not assigned tours. Finally, note that for some $k\in V_{agt}$, $E_{tour}^{k}=\{e:e\in E,\;y_{e,k}=1\}$ for the nonconvex formulation. Hence, for each $e=(i,j)\in E_{tour}^{k}$, we set $y_{e,k}=1$, and set $(z_{e,k,p},z_{e,k,t})=(p_{i,k},t_{i,k})=(z_{e,p},z_{e,t})$ and $(z^{\prime}_{e,k,p},z^{\prime}_{e,k,t})=(p_{j,k},t_{j,k})=(z^{\prime}_{e,p},z^{\prime}_{e,t})$.
@@ IV-C Proof of Validity
The flow constraints in both formulations require $1\leq\alpha\leq m$ vertex-disjoint paths that starts at $s$ and ends at $s^{\prime}$, such that for each target $u\in V_{tar}$, exactly one node in $C_{u}$ is visited by one of the $\alpha$ paths. Consider the edge set $E_{tour}^{k}$ for a feasible path, corresponding to some agent $k\in V_{agt}$. Suppose such a path is chosen by both the formulations. Then, for each $e=(i,j)\in E_{tour}^{k}$, $y_{e,k}=1$ in the nonconvex formulation and $y_{e}=1$ in the MICP. We first show, the cost addend by the two formulations, $l_{e,k}$ and $l_{e}$ are equal for $e\in E_{tour}^{k}$. In the nonconvex formulation, (16) requires $(z_{e,k,p},z_{e,k,t})=(p_{i,k},t_{i,k})$ and $(z^{\prime}_{e,k,p},z^{\prime}_{e,k,t})=(p_{j,k},t_{j,k})$. Additionally, (6) and (12) require $(p_{i,k},t_{i,k})\in X_{i}$ and $(p_{j,k},t_{j,k})\in X_{j}$. These two requirements are the same as saying $(z_{e,k,p},z_{e,k,t})\in X_{i}$, $(z^{\prime}_{e,k,p},z^{\prime}_{e,k,t})\in X_{j}$, and for any two adjacent edges $e=(i,j)$ and $f=(j,k)$ in $E_{tour}^{k}$, $(z^{\prime}_{e,k,p},z^{\prime}_{e,k,t})=(z_{f,k,p},z_{f,k,t})$. Also, in the MICP formulation, constraints (23) to (26) require $(z_{e,p},z_{e,t})\in X_{i}$, $(z^{\prime}_{e,p},z^{\prime}_{e,t})\in X_{j}$. Additionally, the new flow requirement in (22) ensures that for any two adjacent edges $e=(i,j)$ and $f=(j,k)$ in $E_{tour}^{k}$, $(z^{\prime}_{e,p},z^{\prime}_{e,t})=(z_{f,p},z_{f,t})$. Finally, the time-feasibility constraints in both the formulations (13) and (27) have the same form. Hence, for an agent $k\in V_{agt}$, and corresponding edges in $E_{tour}^{k}$, the $z$ variables in both the formulations are restricted to take the same set of feasible values. Now, by letting $(z_{e,k,p},z_{e,k,t})$ and $(z^{\prime}_{e,k,p},z^{\prime}_{e,k,t})$ in the nonconvex formulation to be equal to $(z_{e,p},z_{e,t})$ and $(z^{\prime}_{e,p},z^{\prime}_{e,t})$ respectively in the MICP, the cost addend $l_{e,k}$ defined by constraints (14) and (15) in the nonconvex formulation equals to the cost addend $l_{e}$ defined by constraints (28) and (29) in the MICP formulation. Next, suppose an edge $e\in E$ is not part of any of the agent tours in either of the formulations. Then, for this edge, the cost addend by both the formulations is $l_{e,k}=l_{e}=0$. This is because, in the nonconvex formulation, $y_{e,k}=0$ for all $k\in V_{agt}$ and consequently, constraints (16) require $(z_{e,k,p},z_{e,k,t})=(z^{\prime}_{e,k,p},z^{\prime}_{e,k,t})=(0,0,0)$. Similarly, in the MICP formulation, $y_{e}=0$ and constraints (23) to (26) require $(z_{e,p},z_{e,t})=(z^{\prime}_{e,p},z^{\prime}_{e,t})=(0,0,0)$. A set of agent tours recovered from a feasible solution to the MICP by following the procedure discussed previously, corresponds to a feasible solution for the nonconvex formulation which shares the same objective value as the MICP. Moreover, both the formulations share the same optimal value. Hence, optimal agent tours for the nonconvex formulation can be obtained from an optimal solution to the MICP by following the procedure explained previously.

### v2

@@ Abstract
The Moving-Target Traveling Salesman Problem (MT-TSP) seeks a shortest path for an agent that starts at a stationary depot, visits a set of moving targets exactly once, each within one of their respective time windows, and returns to the depot. In this paper, we introduce a new Mixed-Integer Conic Program (MICP) formulation for the Multi-Agent Moving-Target Traveling Salesman Problem (MA-MT-TSP), a generalization of the MT-TSP involving multiple agents. Our approach begins by restating the current state-of-the-art MICP formulation for MA-MT-TSP as a Nonconvex Mixed-Integer Nonlinear Program (MINLP), followed by a novel reformulation into a new MICP. We present computational results demonstrating that our formulation outperforms the state-of-the-art, achieving up to two orders of magnitude reduction in runtime, and over 90% improvement in optimality gap.
@@ I Introduction
The objective of this paper is to find exact solutions for the Multi-Agent Moving-Target Traveling Salesman Problem (MA-MT-TSP), a generalization of the MT-TSP with multiple agents. Currently, the only exact solver for this problem is the MICP introduced in [27]. Hence, we take this as our baseline. Here, each target is assumed to move along its own line, and have one associated time window. However, in this paper, we consider a less restricted case where each target traverse a path made of piecewise-linear segments, and have more than one associated time window (Fig. 1). Therefore, we first extend the baseline MICP to accommodate these changes. Then, we pose this extended formulation as a nonconvex mixed-integer nonlinear program (nonconvex MINLP), which we then finally reformulate as a new MICP, similar to the works in [23], and [18].
@@ III Baseline MICP for MA-MT-TSP
Although this MICP describes the MA-MT-TSP well, It is challenging to solve in practice. This is possibly due to the big-$M$ constraints (7) and (9), which may lead to poor convex relaxations, as well as numerical instabilities, as noted by the authors in [27]. In the next section, we present an alternative MICP based on the ideas in [23] and [18], that finds the optimum for the MA-MT-TSP, much faster than the baseline. Prior to presenting this formulation, we restate the baseline as a nonconvex MINLP. This will aid us in proving that an optimal solution for our new MICP indeed provides an optimal solution for the MA-MT-TSP.
@@ IV A New MICP for the MA-MT-TSP
IV-A Nonconvex MINLP Formulation for MA-MT-TSP
This section details how the baseline MICP formulation for the MA-MT-TSP can be restated as a nonconvex MINLP. We refer to this program as the nonconvex formulation for simplicity. First, we define the decision variables. For each node $i\in V$ and agent $k\in V_{agt}$, we reuse the variable $t_{i,k}$ from the baseline. In addition, we also define $p_{i,k}$ as a new auxiliary variable. For each edge $e=(i,j)\in E$, we reuse the variables $y_{e,k}$, $l_{e,k}$, and $l_{e,k}^{xy}$ from the baseline. The $l_{e,k}$ variable is defined with the help of new real auxiliary variables $z_{e,k,p}$ and $z_{e,k,t}$ representing the products $y_{e,k}p_{i,k}$ and $y_{e,k}t_{i,k}$ respectively, and variables $z^{\prime}_{e,k,p}$ and $z^{\prime}_{e,k,t}$ representing $y_{e,k}p_{j,k}$ and $y_{e,k}t_{j,k}$ respectively. Observe how in the baseline, $l_{e,k}$ was defined with the help of $\overline{l}_{e,k}$ and big-$M$ constraints instead. We now present the nonconvex formulation below.
@@ IV-A Nonconvex MINLP Formulation for MA-MT-TSP
$\begin{split}&z_{e,k,p}=y_{e,k}p_{i,k},\;z_{e,k,t}=y_{e,k}t_{i,k},\;z^{\prime}_{e,k,p}=y_{e,k}p_{j,k},\\ &z^{\prime}_{e,k,t}=y_{e,k}t_{j,k},\;\;\forall\;e=(i,j)\in E,\;k\in V_{agt}.\end{split}$ (16)
@@ IV-B A New MICP Formulation for MA-MT-TSP
From an optimal solution to the MICP, optimal agent tours for the nonconvex formulation can be obtained by considering the subset of edges $E_{tour}:=\{e\in E:y_{e}=1\}$ and finding all the $1\leq\alpha\leq m$ vertex-disjoint paths starting at $s$ and ending at $s^{\prime}$ within $E_{tour}$. This can be done using a graph search algorithm such as Depth-First Search (DFS). Since all the agents are homogeneous and share the same depot, any of the paths can be assigned to any agent. Let $E_{tour}^{k}\subseteq E_{tour}$ denote the path assigned to an agent $k\in V_{agt}$. Note that $E_{tour}=\bigcup_{k\in V_{agt}}E_{tour}^{k}$. Also note that it is possible $E_{tour}^{k}=\emptyset$ for some $k$ in the case where some of the agents are not assigned tours. Finally, note that for some $k\in V_{agt}$, $E_{tour}^{k}=\{e\in E:\;y_{e,k}=1\}$ for the nonconvex formulation. Hence, for each $e=(i,j)\in E_{tour}^{k}$, we set $y_{e,k}=1$, and set $(z_{e,k,p},z_{e,k,t})=(p_{i,k},t_{i,k})=(z_{e,p},z_{e,t})$ and $(z^{\prime}_{e,k,p},z^{\prime}_{e,k,t})=(p_{j,k},t_{j,k})=(z^{\prime}_{e,p},z^{\prime}_{e,t})$.
@@ IV-C Proof of Validity
Let $E_{tour}$ be the set of all edges selected at optimality by either of the two formulations. For the nonconvex formulation, $E_{tour}:=\{e\in E:\exists\;k\in V_{agt}\;\text{such that}\;y_{e,k}=1\}$, and for the MICP formulation, $E_{tour}:=\{e\in E:y_{e}=1\}$. The flow constraints in both formulations require that the edges in $E_{tour}$ form $1\leq\alpha\leq m$ vertex-disjoint paths, that starts at $s$ and ends at $s^{\prime}$, such that for each target $u\in V_{tar}$, exactly one node in $C_{u}$ is visited by one of the $\alpha$ paths. First, consider an edge $e\notin E_{tour}$. In the nonconvex formulation, for all $k\in V_{agt}$, $y_{e,k}=0$. Consequently, constraints (16) require $(z_{e,k,p},z_{e,k,t})=(z^{\prime}_{e,k,p},z^{\prime}_{e,k,t})=(0,0,0)$ for all $k\in V_{agt}$. Similarly, in the MICP formulation, $y_{e}=0$ leading to constraints (23) to (26) requiring $(z_{e,p},z_{e,t})=(z^{\prime}_{e,p},z^{\prime}_{e,t})=(0,0,0)$. Hence, the cost addend is $\sum_{k\in V_{agt}}l_{e,k}=l_{e}=0$ for either formulation. Now, consider an edge $e=(i,j)\in E_{tour}$. In the nonconvex formulation, since edge $e$ is assigned to some unique agent $k^{*}\in V_{agt}$, $y_{e,k}=1$ when $k=k^{*}$ and $y_{e,k}=0$ for all other $k\in V_{agt}\setminus\{k^{*}\}$. Consequently, (16) requires $(z_{e,k^{*},p},z_{e,k^{*},t})=(p_{i,k^{*}},t_{i,k^{*}})$ and $(z^{\prime}_{e,k^{*},p},z^{\prime}_{e,k^{*},t})=(p_{j,k^{*}},t_{j,k^{*}})$. Additionally, (6) and (12) require $(p_{i,k^{*}},t_{i,k^{*}})\in X_{i}$ and $(p_{j,k^{*}},t_{j,k^{*}})\in X_{j}$. These two requirements, with the help of flow conservation, (5) reads $(z_{e,k^{*},p},z_{e,k^{*},t})\in X_{i}$, $(z^{\prime}_{e,k^{*},p},z^{\prime}_{e,k^{*},t})\in X_{j}$, and for adjacent edges $e=(i,j)$ and $f=(j,k)$ in $E_{tour}$, $(z^{\prime}_{e,k^{*},p},z^{\prime}_{e,k^{*},t})=(z_{f,k^{*},p},z_{f,k^{*},t})$. Similarly, in the MICP formulation, $y_{e}=1$ and therefore, constraints (23) to (26) require $(z_{e,p},z_{e,t})\in X_{i}$, $(z^{\prime}_{e,p},z^{\prime}_{e,t})\in X_{j}$. Additionally, the new requirement in (22) ensures that for adjacent edges $e=(i,j)$ and $f=(j,k)$ in $E_{tour}$, $(z^{\prime}_{e,p},z^{\prime}_{e,t})=(z_{f,p},z_{f,t})$. Finally, the time-feasibility constraints in both the formulations (13) and (27) have the same form. Therefore, corresponding to an edge $e\in E_{tour}$, the cost addend in the nonconvex formulation is $\sum_{k\in V_{agt}}l_{e,k}=l_{e,k^{*}}=\left\lVert z^{\prime}_{e,k^{*},p}-z_{e,k^{*},p}\right\rVert_{2}$ which is the same as the cost addend $l_{e}=\left\lVert z^{\prime}_{e,p}-z_{e,p}\right\rVert_{2}$ by the MICP formulation. Since the $E_{tour}$ corresponding to an optimal solution for one formulation, when used in the other formulation, results in a feasible solution with the same cost, both formulations share the same optimal value. Moreover, from an optimal solution to the MICP, optimal agent tours for the nonconvex formulation can be obtained by following the previously described procedure.